OpenStudy (anonymous):

Find the volume of the solid obtained by rotating about the x-axis the region bounded by the curve y =sqrt(x-1) and the x-axis between x = 2 and x = 5 .

OpenStudy (anonymous):

\[ \int_2^5 \pi y^2 dy \] It should be easy now. Replace y by its value and integrate.

OpenStudy (callisto):

Use disc method~ \[volume = \pi \int_{2}^{5} \sqrt{x-1} ^2 dx =\pi \int_{2}^{5} (x-1)dx \] I bet you can solve it from here :)